ROTATIONAL SPECTRA AND AB INITIO THEORY OF THE SILACUMULENE CHAINS H_{2}C_{n}Si

\underline{MICHAEL C. MCCARTHY}, PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA, 02138, and Division of Engineering and Applied Sciences, 29 Oxford Street, Cambridge, MA, 02138; and PETER BOTSCHWINA, Institut für Physikalische Chemie, Universität Göttingen, Tammannstr. 6, D-37077 Göttingen, Germany.

The silacumulene chains H_{2}C_{n}Si have been the subject of study by molecular beam Fourier transform microwave spectroscopy and ab initio calculation by coupled cluster methods. Each chain is predicted to possess a singlet electronic ground state, and a linear heavyatom backbone with two equivalent off-axis H atoms. Equilibrium structures and dipole moments for chains up to H_{2}C_{8}Si were derived from large-scale CCSD(T) calculations. The rotational spectra of two chains, H_{2}CCSi and H_{2}C_{4}Si, have now been detected in the laboratory, and precise spectroscopic constants were derived for each molecule. For H_{2}CCSi, an experimental structure has been determined to high accuracy by isotopic substitution. In addition, rotational transitions between K_{a} = ±1 levels of the ^{29}Si and the two ^{13}C isotopic species of H_{2}CCSi exhibit nuclear spin-rotation hyperfine structure owing to the abnormally large component of the nuclear spin-rotation tensor along the a-inertial axis.